The Shoe Outlet has paid annual dividends of $0.80, $0.85, $1.05, $1.25, and $1.50 per share over the last five years, respectively. The stock is currently selling for $56.23 a share. What is the cost of equity?

=20.33%

Please check the answers and show all work typed out. No excel or grid style please as I am on mobile.
Answer #1

Growth 1=(.85-.80)/.80=6.25%

Growth 1 =(1.05-.85/.85=23.53%

Growth 1=(1.25-1.05)/1.05=19.05%

Growth 1=(1.50-1.25)/1.25=20.00%

Average Growth =(6.25%+25.53%+19.05%+20%)/4=17.21%

cost of equity=D1/Price +growth

cost of equity=(1.50*(1+17.21%))/56.23+17.21%=20.33%